itsmehjay:

A rectangular pool has a length that is 13 meters longer than its width. Its perimeter is 74 meters. What is the width of this pool? Must include the unit of measurement in your answer.

JamesTDG:

I think your equation is 74 = (13+x)x

AZ:

I disagree

AZ:

Do you know the formula for perimeter? Perimeter = 2 * (Length + width) The length is 13 meters longer than the width If we call the 'width' as 'x' How would you write the length? It's 13 longer than the width

AZ:

longer than just means addition so how would you write that in terms of x and 13 hint: it's addition

JamesTDG:

I kinda forgot to include the '*2' part...

AZ:

and adding the width and length

itsmehjay:

so since i'm adding both of it my equation would look like: 74 = 2(13l + w) ?

AZ:

no

itsmehjay:

;-;

AZ:

the length is not 13 TIMES the width it's 13 MORE than the width so if you want to call your width 'w' how would you describe the length in terms of the width?

JamesTDG:

It is sorta like this: 74 = ((13+x) + x)2

itsmehjay:

74 = 2(13 + w) ?? i have no idea

AZ:

look let's say the width is 2 if the length is 13 MORE THAN the width that means the length is 13 + 2 so instead of saying the width is 2, let's say the width is a variable `w` then what is the length?

IhelpUuHELPme:

no no- the formula for Perimeter is: length + width Width: length - 13 Perimeter is 74 meters: 2(l + w) = 74 l + w = 37

itsmehjay:

15? bc i added

itsmehjay:

eh?

JamesTDG:

13+x = length is 13 MORE THAN the width = length x = width

AZ:

yes but now instead of '2' imagine it's a w can you add 13 + w? you can't until you know what the value of 'w' is so you just write that the length is 13 + w

itsmehjay:

@ihelpuuhelpme wrote:
no no- the formula for Perimeter is: length + width Width: length - 13 Perimeter is 74 meters: 2(l + w) = 74 l + w = 37
how did u get 37?

AZ:

@ihelpuuhelpme wrote:
no no- the formula for Perimeter is: length + width Width: length - 13 Perimeter is 74 meters: 2(l + w) = 74 l + w = 37
is this a sign for me to give up

itsmehjay:

lmaoo no e.e

JamesTDG:

@az wrote:
@ihelpuuhelpme wrote:
no no- the formula for Perimeter is: length + width Width: length - 13 Perimeter is 74 meters: 2(l + w) = 74 l + w = 37
is this a sign for me to give up
Unless you WANT to have to replace your desk...

itsmehjay:

um.. ;-;

AZ:

@az wrote:
look let's say the width is 2 if the length is 13 MORE THAN the width that means the length is 13 + 2 so instead of saying the width is 2, let's say the width is a variable `w` then what is the length?

AZ:

do you know what a variable is?

AZ:

have you ever seen questions using 'x' and you solve for 'x'

itsmehjay:

um yeah duh e.e

AZ:

so essentially, we have TWO variables in this question LENGTH and WIDTH with variables, you can use any letter to represent it so L for Length and W for Width

itsmehjay:

yeahhh..

AZ:

we only have ONE equation though Perimeter = 2 * (Length + Width) when you have ONE equation, you can only solve for ONE variable so that means you have to write the equation all in ONE variable so when they tell you that the length is 13 more than the width that means Length equals to whatever the width is but 13 MORE so Length = Width + 13 does that make sense so far?

itsmehjay:

@az wrote:
we only have ONE equation though Perimeter = 2 * (Length + Width) when you have ONE equation, you can only solve for ONE variable so that means you have to write the equation all in ONE variable so when they tell you that the length is 13 more than the width that means Length equals to whatever the width is but 13 MORE so Length = Width + 13 does that make sense so far?
ohh ya totally e.e

AZ:

okay, you didn't get it so you understand the formula right? perimeter = 2 *(length + width) perimeter is basically adding all the sides up and in a rectangle, you have four sides 

AZ:

the perimeter is going to be equal to length + length + width + width does that much make sense?

itsmehjay:

no dude i told u i understood it lol

AZ:

oh, it sounded sarcastic lol

itsmehjay:

oh sorry .,.

IhelpUuHELPme:

uh- sorry - Formula for Perimeter is: \[2(W) + 2(L)] or \[length + length + width + width\] Width: length - 13 Perimeter is 74 meters: 2(l + w) = 74 So we know that:. - 13 meters LONGER than the width - Perimeter is 74 13m + the width or 13m + w Now, the perimeter equation would be: 2(w) + 2 (w + 13 m) > 2(w) + 2(w) + 2(13m) ⇒ 4(w) + 26m = 74 m. One variable left o-0 subtract 26 from both sides 4(w) which would equal 48m now you have to divide 4 from both sides to find your width width = 12 meters o-o lynn does that make sense? o-0

AZ:

so now your question says perimeter is 74 perimeter = 2 * (length + width) replace perimeter with 74 74 = 2*(length + width) if you want to simplify it, you can just divide by 2 on both sides

AZ:

but then you still won't be able to calculate the length of the width you have to use the other piece of information they gave you length is 13 meters longer than the width so if the width was 2, the length would be 13 + 2 if the width was 100, the length would be 13 + 100 and so on so if the length is equal to the WIDTH + 13 then you can replace length with `width + 13`

AZ:

so 74 = 2 * (length + width) you can replace `length` with `width + 13` and simplify it think of the `width` as `x` 74 = 2 * (x+13 + x)

itsmehjay:

ohh i see then does that mean I have to distribute?

AZ:

you can, or you can first just divide the 2 if you divide by 2, you get 74/2 = (x+13 +x) since the 2 and 2 would cancel out on the right side

itsmehjay:

oh there's 2 2x's so i can subtract one of them to cancel it out the be left with 74/2 = 13?? uh

AZ:

we're not subtracting we're adding x + 13 + x what is an apple + an apple how many apples do you have?

AZ:

also what is 74/2 just divide 74 by 2

itsmehjay:

i got 37 but is that it?

AZ:

no we had 74/2 = (x+13 +x) you said that 74/2 = 37 so now we have 37 = x + 13 + x we need to solve until we get x = ??

itsmehjay:

so do i subtract the 13 to 37 and add the 2 x's to make 2x ?

itsmehjay:

so 37-13 = 2x?

AZ:

yup!

itsmehjay:

YEE

AZ:

Good good!! and what is 37 -13 equal to

JamesTDG:

then divide the left half by 2 to get x

JamesTDG:

(I timed that bad)

itsmehjay:

@az wrote:
Good good!! and what is 37 -13 equal to
24~!

AZ:

now we have 24 = 2x so to solve for 'x' which is the width, you just divide both sides by 2 so x = 24/2 which is equal to what?

itsmehjay:

12 :0

AZ:

Good! And what would the units be? Feet? Meters? Yards? Miles?

itsmehjay:

mEtErS-?

AZ:

You got it!

itsmehjay:

YEEE I feel so accomplished thxx :D

AZ:

you're welcome :)